# Perfect Secret Key Generation for a class of Hypergraphical Sources

### v1

@@ A-A Omitted Proofs of Section III
Let $\mathbf{K}$ denote the vector of random variables corresponding to edges in $\mathcal{E}_{ij}$, and let $\mathbf{U}$ denote the vector of random variables corresponding to edges in $\mathcal{E}_{i}\setminus\mathcal{E}_{ij}$. Since the entries are i.i.d. Be(1/2), the distribution of $\mathbf{K}$ is uniform over $\{0,1\}^{|\mathcal{E}_{ij}|}$. Thus, $H(K^{(i)})=|\mathbf{K}|=\log|\mathcal{K}|$, implying uniformity.
Let $N=|\mathbf{U}|+|\mathbf{K}|=\binom{m}{t}$, be the total dimension. The random variables in $\mathbf{U}$ and $\mathbf{K}$ are i.i.d. Be(1/2) variables, so the joint vector $(\mathbf{U},\mathbf{K})$ is uniformly distributed over $\mathbb{F}_{2}^{N}$. The public communication $F$ consists of a vector of messages $\mathbf{M}$ generated by the rule $\mathbf{M}=\mathbf{U}\oplus G\mathbf{K}$, where $G$ describes the XOR combinations of variables. We analyze the joint distribution of the messages and the key by considering the augmented linear transformation $T:\mathbb{F}_{2}^{N}\to\mathbb{F}_{2}^{N}$:
$\begin{bmatrix}\mathbf{M}\\ \mathbf{K}\end{bmatrix}=\underbrace{\begin{bmatrix}I_{|\mathbf{U}|}&G\\ 0&I_{|\mathbf{K}|}\end{bmatrix}}_{T}\begin{bmatrix}\mathbf{U}\\ \mathbf{K}\end{bmatrix}$.
By construction, $T$ has full rank. Since the input $(\mathbf{U},\mathbf{K})$ is uniform and $T$ is a bijection, the output $(\mathbf{M},\mathbf{K})$ is uniformly distributed over $\mathbb{F}_{2}^{N}$. Hence that $\mathbf{M}$ and $\mathbf{K}$ are independent, i.e., $H(\mathbf{K}\mid\mathbf{M})=H(\mathbf{K})$, yielding $I(\mathbf{K};\mathbf{M})=0$, implying perfect secrecy.
A-B Omitted Proofs of Section IV
@@
Suppose $|\mathcal{M}|=k+1,\ k\geq 4$ (for $k=3$, we simply have a triangle and sending the addition of all messages suffices). Without loss of generality, let the induced cycle source be $\mathcal{H}_{i}=\{[k]\cup\{i\},\mathcal{E}\}$. We introduce the following notation for the edges of the graph: for each $j\in[k-1],e_{j}:=\{i,j,j+1\}$ and $e_{k}=\{i,k,1\}$.
    Case I: $k=3l$ or $k=3l+2$
    The protocol is as follows: the anchor $i$ sends the following k2k-2 messages: $e_{1}\oplus e_{2}\oplus e_{3}$, $e_{2}\oplus e_{3}\oplus e_{4}$, … $e_{k-2}\oplus e_{k-1}\oplus e_{k}$. Once we show recovery, we know that there exists a secret key of at least $2$ bits by Lemma 2.
    To show recovery, first, consider some vertex $j\neq 1$. By virtue of being a member of those edges, vertex $j$ has access to the bits $e_{j-1},e_{j}$. From these, they can recover the bits $e_{j-2},e_{j+1}$ from the communications $e_{j-2}\oplus e_{j-1}\oplus e_{j}$ and $e_{j-1}\oplus e_{j}\oplus e_{j+1}$, respectively. Inductively, they can now recover all the other bits. Now, for vertex $1$, they have access to edges $e_{1}$ and $e_{k}$. By adding the communications $e_{1}\oplus e_{2}\oplus e_{3}$ and $e_{2}\oplus e_{3}\oplus e_{4}$ along with $e_{1}$, they can recover $e_{4}$. In a similar manner, they can recover every bit of the form $e_{3c+1}$.
    If $k=3l$, the last bit vertex $1$ recovers in this sequence is $e_{3l-2}$, which allows them to recover $e_{3l-1}$ from the communication $e_{3l-2}\oplus e_{3l-1}\oplus e_{3l}$. Vertex $1$ now has the same information as vertex $3l$, which we have shown recovers everything, thus, vertex $1$ can recover all the information.
    Similarly, if $k=3l+2$, the last bit recovered in this sequence is $e_{3l+1}$, which allows it to recover all the knowledge of vertex $3l+2$, which suffices again by the same argument as above.
    Case II: $k=3l+1$
    The protocol is as follows: the anchor $i$ sends the messages $e_{1}\oplus e_{2}\oplus e_{3}$, …, $e_{k-3}\oplus e_{k-2}\oplus e_{k-1}$, and finally the message $e_{k-1}\oplus e_{k}\oplus e_{1}$.
    To show recovery, first, consider some vertex $j\notin\{1,k\}$. By virtue of their edges, vertex $j$ has access to $e_{j-1},e_{j}$. From these, they can recover backwards $e_{j-2},\dots,e_{1}$ using the standard communications (e.g., $e_{j-2}\oplus e_{j-1}\oplus e_{j}$). Similarly, they can recover forward $e_{j+1},\dots,e_{k-1}$. Now, possessing both $e_{1}$ and $e_{k-1}$, they can use the last message $e_{k-1}\oplus e_{k}\oplus e_{1}$ to recover $e_{k}$. Thus, they recover all bits.
    For vertex $1$, they have access to edges $e_{1},e_{k}$. From the last message $e_{k-1}\oplus e_{k}\oplus e_{1}$, they immediately recover $e_{k-1}$. As in Case I, by adding communications $e_{1}\oplus e_{2}\oplus e_{3}$ and $e_{2}\oplus e_{3}\oplus e_{4}$ along with $e_{1}$, they can recover $e_{4}$. Continuing this pattern, they recover every bit of the form $e_{3c+1}$. Since $k=3l+1$, we have k3=3l2k-3=3l-2, which is of this form. Thus, vertex $1$ recovers $e_{k-3}$. Now, knowing $e_{k-3}$ and $e_{k-1}$, vertex $1$ uses the message $e_{k-3}\oplus e_{k-2}\oplus e_{k-1}$ to recover $e_{k-2}$. Vertex $1$ now holds $e_{k-2},e_{k-1}$, which recovers the view of party k1k-1 (who we have shown recovers everything).
    For vertex $k$, they have access to $e_{k-1},e_{k}$. Using the last message $e_{k-1}\oplus e_{k}\oplus e_{1}$, they recover $e_{1}$. They now hold $e_{1},e_{k}$, which recovers the view of party $1$. Since party $1$ recovers everything, party $k$ does as well.
@@
We define k=m1k=m-1 and apply the generic scheme as follows:
  1. 1.
    Anchor Selection: We choose the set of anchors to be all vertices $A=[m]$.
  2. 2.
    Projection Analysis: Consider an arbitrary anchor $i\in A$. The projected graph consists of vertices $m]\setminus\{i\}$ and edges $\{e\setminus\{i\}:i\in e\}$. Since the hypergraph is complete, the projected edges connect every pair of vertices in $m]\setminus\{i\}$. Thus, the projected graph is the complete graph $K_{k}$ (where k=m1k=m-1).
We now proceed by analyzing the decomposition and overlap based on the parity of $k$.
Odd Case ($m$ is even)
    Decomposition: Since $k$ is odd, we invoke Walecki’s Theorem [[alspach2008wonderful], Theorem 5.2], which states that the complete undirected graph $K_{k}$ admits a decomposition into $\frac{k-1}{2}$ edge-disjoint Hamiltonian cycles. Thus, we set $p_{i}=\frac{m-2}{2}$ and use a single copy of the projection ($k_{i}=1$).
    Overlap ($n$): Every edge $e\in\mathcal{E}$ contains exactly $3$ vertices. Since $A=[m]$, all three vertices act as anchors for the edge. With $k_{i}=1$, the total usage count for each edge is $t_{e}=3\times 1=3$. We therefore set $n=3$.
Using the rate formula derived in the generic scheme, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}$. Substituting $|A|=m$, $p_{i}=(m-2)/2$, and $n=3$ gives us $R=\frac{2\cdot m\cdot\frac{m-2}{2}}{3}=\frac{m(m-2)}{3}$. This achieves capacity, as the singleton bound for the complete hypergraph is $\frac{2}{m-1}\binom{m}{3}=\frac{m(m-2)}{3}$.
Even Case ($m$ is odd)
For $k\in\{4,6\}$, we provide explicit decompositions in Figures 2 and 2. For other values:
    Decomposition: Since $k$ is even, we consider the complete directed graph $\overrightarrow{K}_{k}$. Tillson’s Theorem [Tillson1980AHD] states that for $k\notin\{4,6\}$, $\overrightarrow{K}_{k}$ admits a decomposition into k1k-1 edge-disjoint directed Hamiltonian cycles. To utilize this, we treat $2$ copies of the undirected projection as a single complete directed graph. Thus, we set $k_{i}=2$ and obtain a decomposition size of $p_{i}=m-2$.
    Overlap ($n$): Every edge contains 3 anchors. Since we use $k_{i}=2$ copies for every anchor, each edge is counted $3\times 2=6$ times across the projections. We therefore set $n=6$.
Using the rate formula, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}$. Substituting $|A|=m$, $p_{i}=m-2$, and $n=6$ gives us $R=\frac{2\cdot m\cdot(m-2)}{6}=\frac{m(m-2)}{3}$. This rate matches the singleton bound capacity of $\frac{m(m-2)}{3}$.
[…]
(a) Copy 1
[…]
(b) Mixed (Copy 1 + Copy 2)
[…]
(c) Copy 2
Figure 1: Hamiltonian cycle packing for $K_{4}$.
[…]
(d) Copy 1
[…]
(e) Copy 1
[…]
(f) Mixed (uses $1\!\to\!2,\,3\!\to\!6$)
[…]
(g) Mixed (uses $4\!\to\!5$)
[…]
(h) Copy 2
Figure 2: Hamiltonian cycle packing for $K_{6}$.
@@
We apply the generic scheme as follows:
  1. 1.
    Anchor Selection: We choose the singleton set of anchors $A=\{i\}$.
  2. 2.
    Projection Analysis: The projected graph at vertex $i$ consists of vertices $\mathcal{M}\setminus\{i\}$ and edges $\{e\setminus\{i\}:e\in\mathcal{E}\}$.
  3. 3.
    Decomposition: We assume that $k$ independent copies of this projected graph admit a Hamiltonian cycle packing of size $p$. Thus, we have a total decomposition size of $p$ distributed over these copies.
  4. 4.
    Overlap ($n$): We set the number of copies of the original source to $n=k$. Since $i$ is in every edge and we consider $k$ copies of the projection, the overlap aligns with the decomposition parameters.
Using the rate formula derived in the generic scheme, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}.$ Substituting the parameters $p_{i}=p$ and $n=k$ gives us $R=\frac{2p}{k}.$ We now verify that this achieves capacity. First, observe that a single Hamiltonian cycle on the projected graph (which has m1m-1 vertices) contains m1m-1 edges. Therefore, the union of $p$ disjoint cycles contains p(m1)p(m-1) edges. Since these edges are distributed over $k$ copies, the number of edges in a single copy of the hypergraph is $|\mathcal{E}|=\frac{p(m-1)}{k}.$ The multi-information for the singleton partition $X_{S}$, denoted $I(X_{S})$, provides an upper bound on the secret key capacity as I(X_S) = 2m-1—E— = 2m-1 ( p(m-1)k ) = 2pk. Since our achievable rate matches the upper bound derived from the singleton partition, the singleton partition must minimize the expression for capacity. This demonstrates that the source is Type S.
@@
  1. 1.
    Anchor Selection: We choose the set of anchors to be $A=V_{1}$.
  2. 2.
    Projection Analysis: Consider an arbitrary anchor $i\in A$. The projected graph consists of vertices $V\setminus\{i\}$ and edges $\{e\setminus\{i\}:i\in e\}$. Since every edge in $\mathcal{E}$ containing $i$ must have its second vertex in $V_{1}\setminus\{i\}$ and its third vertex in $V_{2}$, the projected edges connect every node in $V_{1}\setminus\{i\}$ to every node in $V_{2}$. Observing that $|V_{1}\setminus\{i\}|=2r$ and $|V_{2}|=2r$, the projected graph is the complete bipartite graph $K_{2r,2r}$.
  3. 3.
    Decomposition: It is a known result that the complete bipartite graph $K_{2n,2n}$ admits a decomposition into $n$ Hamiltonian cycles. Thus, for our projected graph $K_{2r,2r}$, we have a Hamiltonian decomposition of size $p_{i}=r$. We set $k_{i}=1$ (a single copy).
  4. 4.
    Overlap ($n$): Every edge $e\in\mathcal{E}$ is of the form $\{u,v,w\}$ with $u,v\in V_{1}$ and $w\in V_{2}$. Since our anchor set is $A=V_{1}$, the anchors present in edge $e$ are exactly $u$ and $v$. Consequently, every edge is used exactly twice, implying $n=2$.
Using the rate formula derived in the generic scheme, we have $R=\frac{2\sum_{i\in A}p_{i}}{n}$. Substituting our parameters $|A|=2r+1$, $p_{i}=r$, and $n=2$ gives us $R=\frac{2\cdot(2r+1)\cdot r}{2}=r(2r+1).$ We now verify that this achieves capacity. The total number of edges in the graph is $|\mathcal{E}|=\binom{2r+1}{2}\times 2r=r(2r+1)(2r)$. The multi-information for the singleton partition $X_{S}$, denoted $I(X_{S})$, provides an upper bound on the secret key capacity as: $I(X_{S})=\frac{2}{m-1}|\mathcal{E}|=\frac{2}{4r}\left(2r^{2}(2r+1)\right)=r(2r+1)$. Since our achievable rate matches the upper bound derived from the singleton partition, the singleton partition minimizes the expression for capacity. Thus, this source is Type S.

### v2

@@ Appendix C Omitted Proofs of Section IV
For $m=4$, we simply have the anchor send the XOR of all the bits associated with all three hyperedges. Then, every party, having access to at least two hyperedges, recovers the random bits associated with all the hyperedges. Hence, we turn our attention to the case where $m=k+1,k\geq 4$.
We only need to show the necessary m3m-3 bit linear communication achieving perfect omniscience. The 2-bit perfect secret key then follows immediately from Lemma II-B by noting $|\mathcal{E}|=m-1$ as $\mathcal{H}$ is cycle inducing. To define the linear communication, without loss of generality, let party $k+1$ be the anchor, and let the set of edges of $P_{\mathcal{H},k+1}$ be $\{\{j,j+1\}:1\leq j\leq k-1\}\cup\{\{1,k\}\}$. We introduce the following notation for the hyperedges in $\mathcal{E}$: for each $1\leq j\leq k-1,e_{j}:=\{k+1,j,j+1\}$ and $e_{k}=\{k+1,k,1\}$.
Case I: $k=3l$ or $k=3l+2$
The protocol is as follows: the anchor $k+1$ sends the following k2k-2 messages: $\xi_{e_{1}}\oplus\xi_{e_{2}}\oplus\xi_{e_{3}}$, $\xi_{e_{2}}\oplus\xi_{e_{3}}\oplus\xi_{e_{4}},\ldots,\xi_{e_{k-2}}\oplus\xi_{e_{k-1}}\oplus\xi_{e_{k}}$. The number of bits of messages thus sent is k2=m3k-2=m-3.
To show perfect omniscience, first, consider some party $j\neq 1$. By virtue of being a member of those hyperedges, party $j$ has access to the bits $\xi_{e_{j-1}},\xi_{e_{j}}$. From these, they can recover the bits $\xi_{e_{j-2}},\xi_{e_{j+1}}$ from the communications $\xi_{e_{j-2}}\oplus\xi_{e_{j-1}}\oplus\xi_{e_{j}}$ and $\xi_{e_{j-1}}\oplus\xi_{e_{j}}\oplus\xi_{e_{j+1}}$, respectively. Inductively, they can now recover all the other bits. Now, for party $1$, they have access to edges $\xi_{e_{1}}$ and $\xi_{e_{k}}$. By adding the communications $\xi_{e_{1}}\oplus\xi_{e_{2}}\oplus\xi_{e_{3}}$ and $\xi_{e_{2}}\oplus\xi_{e_{3}}\oplus\xi_{e_{4}}$ along with the knowledge of $\xi_{e_{1}}$, they can recover $\xi_{e_{4}}$. Inductively, they can recover every bit of the form $\xi_{e_{3c+1}}$.
If $k=3l$, the last bit party $1$ recovers in this sequence is $\xi_{e_{3l-2}}$, which allows them to recover $\xi_{e_{3l-1}}$ from the communication $\xi_{e_{3l-2}}\oplus\xi_{e_{3l-1}}\oplus\xi_{e_{3l}}$. Party $1$ now has the same information as party $3l$, which we have shown to be able to recover everything. Therefore, party $1$ can recover all the information.
Similarly, if $k=3l+2$, the last bit recovered in this sequence is $\xi_{e_{3l+1}}$, which allows party $1$ to recover all the knowledge of party $3l+2$, which suffices again by the same argument as above.
Case II: $k=3l+1$
The protocol is as follows: the anchor $k+1$ sends the messages $\xi_{e_{1}}\oplus\xi_{e_{2}}\oplus\xi_{e_{3}},\ldots,\xi_{e_{k-3}}\oplus\xi_{e_{k-2}}\oplus\xi_{e_{k-1}}$, and finally the message $\xi_{e_{k-1}}\oplus\xi_{e_{k}}\oplus\xi_{e_{1}}$. The number of bits of messages thus sent is k2=m3k-2=m-3.
To show recovery, first, consider some party $j\notin\{1,k\}$. Party $j$ has access to $\xi_{e_{j-1}},\xi_{e_{j}}$. Using these and the communication, an argument exactly same as that in the previous case shows that they can recover $\xi_{e_{1}},\ldots,\xi_{e_{k-1}}$. Now, possessing both $\xi_{e_{1}}$ and $\xi_{e_{k-1}}$, they can use the last message $\xi_{e_{k-1}}\oplus\xi_{e_{k}}\oplus\xi_{e_{1}}$ to recover $\xi_{e_{k}}$. Thus, they recover all bits.
Party $1$, having access to $\xi_{e_{1}},\xi_{e_{k}}$, immediately recovers $\xi_{e_{k-1}}$ from the last message $\xi_{e_{k-1}}\oplus\xi_{e_{k}}\oplus\xi_{e_{1}}$. As in Case I, party $1$ also recovers bits associated with hyperedges of the form $\xi_{e_{3c+1}}$. Since $k=3l+1$, and so k3=3(l1)+1k-3=3(l-1)+1, the last bit party $1$ recovers in this way is $\xi_{e_{k-3}}$. Now, using $\xi_{e_{k-3}}$ and $\xi_{e_{k-1}}$, party $1$ uses the message $\xi_{e_{k-3}}\oplus\xi_{e_{k-2}}\oplus\xi_{e_{k-1}}$ to recover $\xi_{e_{k-2}}$. Party $1$ thus recovers the view of party k1k-1, which has already been shown to recover all the bits.
For party $k$, we show that it can recover the view of party $1$, which has already been shown to recover everything. Using the last message $\xi_{e_{k-1}}\oplus\xi_{e_{k}}\oplus\xi_{e_{1}}$, they recover $\xi_{e_{1}}$. Together with $\xi_{e_{k}}$, this recovers the view of party $1$ as required. ∎
@@
Set $\delta:=\sum_{i\in A}p_{i}$. We begin with the claim that for blocklength $n$ given in the hypothesis of Theorem 9, there exists a packing of $\mathcal{H}$ into $\mathcal{E}_{(1)},\ldots,\mathcal{E}_{(\delta)}\subseteq\mathcal{E}^{n}$ such that each of the hypergraphs $\mathcal{H}_{(l)},1\leq l\leq\delta,$ are cycle inducing. Therefore, by Proposition 4, each $\mathcal{H}_{(l)}$ generates 2 bits of perfect secret key. This leads to a perfect secret key of rate $\frac{2}{n}\sum_{i\in A}p_{i}$ by using the Lemma II-B.
To complete the proof, it remains to prove the claim that the required packing exists. To facilitate the proof, we first order the set $A$ as $A=\{u_{1},u_{2},\ldots,u_{|A|}\}$, with $u_{1}<u_{2}<\ldots<u_{|A|}$. Next, we reindex the sets $\mathcal{E}_{(l)}$ as $\mathcal{E}_{(i,j_{i})},1\leq i\leq|A|,1\leq j_{i}\leq p_{u_{i}}$. Fix any $e\in\mathcal{E}$, and define $A_{e}=\{i\in A:i\in e\}$, and let $A_{e}=\{u_{l_{1}},u_{l_{2}},\ldots,u_{l_{|A_{e}|}}\}$, where $l_{1}<l_{2}<\ldots<l_{|A_{e}|}$. Next, for any $l\in A$, let $C_{l}^{(1)},\ldots,C_{l}^{(p_{l})}$ denote the $p_{l}$ cycles obtained by the Hamiltonian packing of $k_{l}$ copies of $P_{\mathcal{H},l}$. We say that $e$ takes part in the cycle $C_{l}^{s}$ if $e=\{l\}\cup e^{\prime}$, where $e^{\prime}$ is some edge of the cycle $C_{l}^{(s)}$.
We build the sets $\mathcal{E}_{(i,j_{i})}$ as follows. For any $1\leq q\leq|A_{e}|,$ and any $1\leq j_{l_{q}}\leq p_{u_{l_{q}}}$, assign the fictitious copy $e(\sum_{r=1}^{q-1}k_{l_{r}}+j_{l_{q}})$ to $\mathcal{E}_{(l_{q},j_{l_{q}})}$ provided $e$ takes part in $C_{u_{l_{q}}}^{(j_{l_{q}})}$. Then, by construction, the sets $\mathcal{E}_{(i,j_{i})}$ are disjoint. Furthermore, we claim that $e$ never runs out of fictitious copies needed for the assignment process described above. To see this, note that for any $l_{q},1\leq q\leq|A_{e}|$, the hyperedge $e$ takes part in at most $k_{u_{l_{q}}}$ of the cycles $C_{u_{l_{q}}}^{(s)},1\leq s\leq p_{u_{l_{q}}}$. Thus, the maximum number of copies of $e$ required by the assignment process is $\sum_{q=1}^{|A_{e}|}k_{l_{q}}=\sum_{l\in A}k_{l}\mathbf{1}\{l\in e\}=t_{e}\leq n$. Hence, $e$ never runs out of fictitious copies. Finally, observe that the assignment process ensures that $\mathcal{E}_{(i,j_{i})}$ gets exactly one fictitious copy of every $e$ taking part in $C_{u_{i}}^{(j_{i})},1\leq j_{i}\leq p_{u_{i}},1\leq i\leq|A|$. Hence, with a slight abuse of notation, $\mathcal{H}_{(i,j_{i})}$ is cycle inducing with anchor at $u_{i}$, as required. This completes the proof of the claim, and hence, the proof of Theorem 9. ∎
[…]
(a) Cycle 1
[…]
(b) Cycle 2
[…]
(c) Cycle 3
Figure 1: Hamiltonian cycle packing for $K_{4}$. The edges originating from the two distinct copies are respectively coloured red and blue.
[…]
(d) Cycle 1
[…]
(e) Cycle 2
[…]
(f) Cycle 3
[…]
(g) Cycle 4
[…]
(h) Cycle 5
Figure 2: Hamiltonian cycle packing for $K_{6}$. The edges originating from the two distinct copies are respectively coloured red and blue.
@@
Observe that $A=\mathcal{M}$, and let q=m1q=m-1. We break the proof into the following pair of cases.
Case 1: $m$ is even, that is $q$ is odd.
In this case, we show that for every $i\in\mathcal{M}$, a single copy of $P_{K_{m,3},i}$ admits a Hamiltonian decomposition into $\frac{q-1}{2}$ cycles. Hence, the edge reuse number for every hyperedge is $t_{e}=\sum_{i\in\mathcal{M}}\mathbf{1}\{i\in e\}=3$. Thus, using a blocklength of $n=3$, Theorem 9 ensures that a perfect secret key of rate $\frac{2}{3}\sum_{i\in\mathcal{M}}\frac{q-1}{2}=\frac{m(m-2)}{3}$ can be generated. Lemma 1 thus implies that this perfect secret key achieves capacity.
It remains to show that $P_{K_{m,3},i}$ admits a Hamiltonian decomposition into $\frac{q-1}{2}$ cycles. Firstly, observe that $P_{K_{m,3},i}$ is nothing but the complete graph on m1=qm-1=q vertices. Since $q$ is odd, we invoke Walecki’s Theorem [alspach2008wonderful], Theorem 5.2, which states that the complete graph on an odd number (say $q$) of vertices admits a Hamiltonian decomposition of size $\frac{q-1}{2}$. This completes the proof of Case 1.
Case 2: $m$ is odd, that is $q$ is even.
In this case, we show that for every $i\in\mathcal{M}$, two copies of $P_{K_{m,3},i}$ admit a Hamiltonian decomposition into q1q-1 cycles. Hence, the edge reuse number for every hyperedge is $t_{e}=\sum_{i\in\mathcal{M}}2\mathbf{1}\{i\in e\}=6$. Thus, using a blocklength of $n=6$, Theorem 9 ensures that a perfect secret key of rate $\frac{2}{6}\sum_{i\in\mathcal{M}}(q-1)=\frac{m(m-2)}{3}$ can be generated. Lemma 1 thus implies that this perfect secret key achieves capacity.
It remains to show that two copies of $P_{K_{m,3},i}$ admit a Hamiltonian decomposition into q1q-1 cycles. Recall that $P_{K_{m_{3}},i}$ is essentially the complete graph $K_{q}$ on m1=qm-1=q vertices. For $q\in\{4,6\}$, we provide the explicit decompositions in Figures 2 and 2. For $q\geq 8$, we make use of a result by Tillson [Tillson1980AHD], which states that the complete directed hypergraph on $q$ admits a decomposition into q1q-1 directed Hamiltonian cycles, provided $q\geq 8$. Now, a complete directed graph can be viewed as two copies of the complete graph with the orientations of the edges being discarded. Thus, Tillson’s theorem provides a Hamiltonian decomposition of two copies of $K_{q}$ into q1q-1 Hamiltonian cycles as required. ∎
@@
Choose $A=\{i\}$, and recall that the hypothesis of Corollary 5 states $P_{\mathcal{H},i}$ admits a Hamiltonian decomposition into say $p$ cycles. Thus, edge reuse number $t_{e}=1$ for every $e\in\mathcal{E}$, and hence choose $n=1$. Therefore, the scheme of Theorem 9 generates a perfect secret key of rate $2p$.
Now, $P_{\mathcal{H},i}$ being a graph on m1m-1 vertices, each of the Hamiltonian cycles contains m1m-1 edges. Next, note that if a graph admits a Hamiltonian decomposition, then each edge must participate in exactly one of the cycles. Therefore, we must have $|\mathcal{E}|=|\mathcal{E}_{i}|=p(m-1)$. Therefore, we have by (3) that C_SK(X_M)≤I_S(X_M)=2m-1p(m-1)=2p≤C_SK(X_M), where the last inequality follows by noting that Theorem 9 ensures a key of rate $2p$. Thus, $C_{\mathrm{SK}}(X_{\mathcal{M}})=I_{\mathcal{S}}(X_{\mathcal{M}})=2p$, and hence the source is Type-$\mathcal{S}$. Also, the perfect secret keys generated by Theorem 9 achieve capacity.
@@
Choose $A=V_{1}$. We first claim that for any $i\in V_{1}$, the graph $P_{\mathcal{H},i}$ is the complete bipartite graph $K_{2r,2r}$. This follows by noting every edge in $P_{\mathcal{H},i}$ is of the form $\{j_{1},j_{2}\}$ where $j_{1}\in V_{1}\setminus\{i\}$ and $j_{2}\in V_{2}$. Hence, $P_{\mathcal{H},i}$ is bipartite with two parts $V_{1}\setminus\{i\}$ and $V_{2}$, and $|V_{1}\setminus\{i\}|=|V_{2}|=2r$. Furthermore, every subset $\{j_{1},j_{2}\}$ with $j_{1}\in V_{1}\setminus\{i\}$ and $j_{2}\in V_{2}$ is an edge in $P_{\mathcal{H},i}$, thanks to the hyperedge $\{i,j_{1},j_{2}\}\in\mathcal{E}$. Thus, $P_{\mathcal{H},i}$ is indeed $K_{2r,2r}$.
It is a standard result of graph theory that a single copy of $K_{2r,2r}$ admits a Hamiltonian decomposition of size $r$ – see, for example, the Introduction section of [LA]. Thus, for every $i\in A$, we have $k_{i}=1$ and $p_{i}=r$. Furthermore, note that for every $e\in\mathcal{E}$ the edge reuse number $t_{e}=2$. This follows by noting that any hyperedge $\{j_{1},j_{2},j_{3}\}$, with $j_{1},j_{2}\in V_{1}$ and $j_{3}\in V_{2}$, is used twice, once in $P_{\mathcal{H},j_{1}}$ and once in $P_{\mathcal{H},j_{2}}$. Then, invoking the scheme of Theorem 9 with a blocklength of $n=2$ gives us a perfect secret key of rate $\sum_{i\in A}r=r(2r+1)$.
Finally, we have using (3) that C_SK(X_M)≤I_S(X_M)=24r—E—=12r2r(2r+1)22r=r(2r+1)≤C_SK(X_M), where the last inequality follows by noting that Theorem 9 gives a perfect secret key of rate $r(2r+1)$. Therefore, this perfect secret key achieves capacity, and the source is Type-$\mathcal{S}$.